OpenStudy (anonymous):

Suppose xy=−3 and dt/dy=−1. Find dt/dx when x=2.

OpenStudy (stamp):

\[xy=-3,\ dt/dy=-1\]Find dt/dx when x = 2. \[dt/dx=dy/dx*dt/dy\]Find dy/dx.\[y=-3/x\]\[dy/dx=-3(-1/x^2)\]\[dy/dx=3/x^2\]Substitute dy/dx and dt/dy to find dt/dx.\[dt/dx=(3/x^2)(-1)\]\[t'(x)=-3/x^2\]\[t'(2)=-3/(2)^2\]\[t'(2)=-3/4\]

OpenStudy (anonymous):

it looks right i checked it, but the website i put it on doesnt accept it. is there any other steps? it doesnt seem there shud be